### v1

# MentalMAC: Enhancing Large Language Models for Detecting Mental Manipulation via Multi-Task Anti-Curriculum Distillation

Abstract
Mental manipulation is a subtle yet pervasive form of psychological abuse that poses serious threats to mental health. Its covert nature and the complexity of manipulation strategies make it challenging to detect, even for state-of-the-art large language models (LLMs). This concealment also hinders the manual collection of large-scale, high-quality annotations essential for training effective models. Although recent efforts have sought to improve LLM’s performance on this task, progress remains limited due to the scarcity of real-world annotated datasets. To address these challenges, we propose MentalMAC, a multi-task anti-curriculum distillation method that enhances LLMs’ ability to detect mental manipulation in multi-turn dialogue. Our approach includes: (i) EvoSA, an unsupervised data expansion method based on evolutionary operations and speech act theory; (ii) teacher-model-generated multi-task supervision; and (iii) progressive knowledge distillation from complex to simpler tasks. We then constructed the ReaMent dataset with 5,000 real-world dialogue samples, using a MentalMAC-distilled model to assist human annotation. Vast experiments demonstrate that our method significantly narrows the gap between student and teacher models and outperforms competitive LLMs across key evaluation metrics. All code, datasets, and checkpoints will be released upon paper acceptance.
Warning: This paper contains content that may be offensive to readers.
1 Introduction
Figure 1: A dialogue that involves mental manipulation but remains undetected by both GPT-4 with intent-aware prompting (Ma et al., 2025) and DeepSeek-R1 (Guo et al., 2025). Person1 (left) is the target; Person2 (right) is the manipulator.
Mental manipulation is a subtle and covert form of psychological control that employs language to influence and alter another person’s perception for the manipulator’s own benefit. This type of manipulation is widespread in interpersonal relationships and can have serious negative impacts on mental health (Wang et al., 2024). On social networks, perpetrators often use deception and intimidation to coerce children into engaging in harmful online behaviors, which can lead to long-lasting psychological consequences such as anxiety, depression, and self-harm (Ramiro et al., 2019; World Health Organization, 2024). Within intimate relationships, manipulative tactics like gaslighting systematically distort victims’ sense of reality, resulting in emotional instability and an increased risk of suicidal thoughts (Lohmann et al., 2024; Munro et al., 2024). Furthermore, as AI chatbots become increasingly sophisticated and interactive, new cases of mental manipulation have arisen, with reports highlighting the induction of dependency in adults and the troubling normalization of violent ideologies among children (National Public Radio, 2024; Laestadius et al., 2024).
Unlike other harmful content, mental manipulation is often concealed within a multi-turn dialogue, making it difficult to detect even for experts (Myhill and Hohl, 2019). Wang et al. (2024) introduced the MentalManip dataset and found manipulative dialogues are semantically similar to non-manipulative ones, highlighting the challenge of detection based solely on lexical or semantic cues. Although large language models (LLMs) have advanced many tasks (Yuan et al., 2024), several state-of-the-art (SOTA) LLMs still perform poorly on this task (see Figure 1). First, data resources are limited and difficult to expand due to the covert nature of manipulation; despite LLMs’ strong rewriting abilities, manual annotation remains necessary due to their misalignment with human cognition (Wang et al., 2024). Second, LLMs struggle to understand deep semantics and speech acts in conversations, often confusing manipulative with non-manipulative interactions involving persuasion or arguments. While chain-of-thought (CoT) (Yang et al., 2024) and intent-aware prompting (Ma et al., 2025) provide some improvement, gains remain minimal. Third, the sole dataset for this task, MentalManip, is derived from the Cornell Movie Dialogs Corpus [^1] https://www.cs.cornell.edu/~cristian/Cornell_Movie-Dialogs_Corpus.html and contains only 4,000 samples. The scarcity of real-world dialogue datasets significantly limits progress.
Figure 2: Overall workflow of the proposed MentalMAC. In Step 1, we introduce EvoSA, a novel prompting method designed for the unsupervised generation of dialogues containing or not containing mental manipulation. In Step 2, we employ a teacher model to generate both correct and incorrect rationales, along with feedback for the incorrect ones, to construct multi-task training data. In Step 3, we apply anti-curriculum distillation, training the student model from hard to easy tasks to enhance its performance.
To address the limitations above, we propose MentalMAC [^2] Mental manipulation detection via Multi-task Anti-Curriculum distillation, a novel approach that enhances LLMs’ ability to detect mental manipulation through multi-task anti-curriculum distillation. Figure 2 illustrates the overall workflow. Since mental manipulation is pervasive yet subtle, reliable data are scarce. To tackle this, we introduce EvoSA [^3] A prompting method based on Evolutionary operations and Speech Act theory, enabling unsupervised data expansion. To deepen the student model’s grasp of semantic and manipulative cues, we distill knowledge after data expansion as follows: the teacher model generates rationales for both correct and incorrect answers, with feedback highlighting errors in the latter. From this, we define three tasks: (1) providing feedback on incorrect rationales, (2) judging and justifying dialogues, and (3) binary classification. To better align with cognitive demands, MentalMAC structures distillation in order from hard to easy. It starts with feedback on incorrect rationales, then combines feedback and judgment, and finally integrates all three tasks. This sequencing guides learning in a goal-oriented and structured way, which we call anti-curriculum distillation.
Given the limited data in this domain, we propose ReaMent [^4] Real-world Mental manipulation, a dataset of 5,000 real-world dialogues curated from YTD-18M (Han et al., 2023) for mental manipulation detection. To reduce manual labor, we leveraged a student model distilled via MentalMAC to perform initial screening over a large set of dialogues, identifying those that potentially contain elements of mental manipulation. Human annotators then reviewed and corrected the results to ensure label accuracy. In contrast to MentalManip, which uses scripted movie dialogues, ReaMent offers a more realistic representation of mental manipulation in natural settings. Our key contributions are summarized as follows:
  • MentalMAC enhances LLMs’ ability to detect mental manipulation through three stages: First, EvoSA enables unsupervised data expansion by teachers; Second, the teacher generates multi-task data; Third, tasks are arranged from hard to easy to enable anti-curriculum distillation for student models.
  • ReaMent is a dataset of 5,000 real-world dialogues, constructed with human annotation assisted by a student model distilled via MentalMAC for mental manipulation detection.
  • Vast experiments show that our method not only narrows the gap between student and teacher models but also outperforms SOTA LLMs on several key metrics.
2 Related Works
2.1 Mental Manipulation Detection
Mental manipulation has recently drawn increased attention due to its subtle and insidious nature. However, only one dataset, MentalManip, exists for this task, containing just 4,000 samples. This limited resource highlights the urgent need for larger, high-quality datasets, motivating the construction of ReaMent. Current detection methods mostly rely on large LLMs or commercial APIs with prompting strategies such as few-shot learning and chain-of-thought, which incur high costs (Yang et al., 2024). Although Ma et al. (2025) proposed intent-aware prompting to improve LLMs’ understanding of dialogue intent, improvements remain limited. The scarcity of targeted research reveals a significant gap in this area. To address it, we propose MentalMAC, a method that distills the capabilities of large teacher models into smaller student models, enhancing student performance and potentially enabling it to surpass the teacher.
2.2 Knowledge Distillation for LLMs
Knowledge distillation (KD) (Hinton et al., 2015) is a common technique for transferring knowledge—from a larger teacher model to a smaller student—using hard or soft labels. With Chain-of-Thought (CoT) prompting enhancing LLM reasoning, studies like PINTO (Wang and Chang, 2022), Distilling Step-by-Step (Hsieh et al., 2023), RevThink (Chen et al., 2024), and SuperCorrect (Yang et al., 2025a) have explored multi-task fine-tuning that leverages reasoning explanations and LLM-generated task content. For toxic text detection, Zhang et al. (2024) proposed BD-LLM, which improves detection via bootstrapped prompting and rationale-based distillation. However, BD-LLM depends solely on predefined toxicity definitions and uses a simple rationale distillation, limiting its adaptation to the covert, context-sensitive nature of mental manipulation detection. In contrast, MentalMAC is tailored for detecting mental manipulation in multi-turn dialogues, requiring the student model to master not only correct reasoning but also the identification of errors in deceptive rationales. Inspired by anti-curriculum learning (Wu et al., 2021), we design a progressive distillation schedule that guides the student model from more complex feedback tasks to simpler binary classification, thereby better aligning the training process with the intricate cognitive demands of this task.
3 Method
We propose MentalMAC (Figure 2), which begins with unsupervised augmentation of limited training data, where the teacher generates both correct and incorrect rationales with corresponding answers and feedback identifying errors. These enriched samples are then structured into three increasingly simpler tasks for staged distillation.
3.1 EvoSA-Based Data Expansion
To address the challenge of lack of sufficient training data, we propose EvoSA, inspired by evolutionary operations (Yang et al., 2025b) and speech act theory (Searle et al., 1980), to drive LLMs to unsupervised data augmentation.
Figure 3: EvoSA prompts for dialogue generation. To enable unsupervised dialogue synthesis, Dialogue 1 and Dialogue 2 are randomly selected with the same label as parent dialogues. An initial child dialogue is generated using selection, crossover, and mutation (orange text guides the LLM to focus on speech acts and dialogue elements). The LLM then polishes the child dialogue, analyzes the presence or absence of mental manipulation in the parent dialogues, and revises the child dialogue to ensure label consistency (blue text reflects the parent dialogues label).
Evolutionary algorithms rely on selection, crossover, and mutation to maintain diversity and adaptability within populations of generated individuals. Besides, speech act theory reveals how the intentions behind our words shape communication in varied contexts, reflecting the rich diversity of conversational goals. Drawing inspiration from both, we introduce EvoSA (Figure 3). The process begins by randomly sampling training dialogues that share the same label, which we call parent dialogues. Guided by evolutionary operations and grounded in speech act theory, LLMs then create novel, distinct child dialogues through mutation and recombination (steps 1–3). To guarantee the quality and consistency of labels, the LLM is tasked with polishing the child dialogue, analyzing the reasoning behind the parent label, and revising the child dialogue accordingly to ensure it remains aligned with the original intent (steps 4–7).
Mathematically, we have a pre-trained teacher model $\mathcal{T}$. Let $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{n}$ denote the training dataset, where $x_{i}$ represents the $i$-th dialogue, $y_{i}$ is the corresponding binary label indicating the presence or absence of mental manipulation, and $n$ is the total number of training examples. The unsupervised generation of new data by the teacher model under the guidance of EvoSA prompts can be formalized as
$(x_{k}^{\prime},y_{k}^{\prime})=$ $\mathcal{T}((x_{i},y_{i}),(x_{j},y_{j}),p_{EvoSA})$ (1)
[…]
where $(x_{k}^{\prime},y_{k}^{\prime})$ denotes the newly generated data, and $p_{EvoSA}$ represents our designed EvoSA prompts. It is useful to assume that the original training set $\mathcal{D}$ contains $n_{+}$ samples with mental manipulation and $n_{-}$ samples without it, where $n_{+}+n_{-}=n$. Then the expanded training dataset $\mathcal{D^{\prime}}$ can be denoted as
$\mathcal{D^{\prime}}=\mathcal{D}$$\cup\{(x_{i}^{\prime+},y_{i}^{\prime+})\}_{i=1}^{n_{+}^{\prime}}\cup\{(x_{j}^{\prime-},y_{j}^{\prime-})\}_{j=1}^{n_{-}^{\prime}}$ (2)
[…]
$\;\;\;\;\;\;\;\;n_{-}^{\prime}\in[1,C_{n_{-}}^{2}]\cap\mathbb{Z}$
where $\{(x_{i}^{\prime+},y_{i}^{\prime+})\}$ and $\{(x_{j}^{\prime-},y_{j}^{\prime-})\}$ denote new generated data labelled as containing and not containing mental manipulation, respectively. Besides, $n_{+}^{\prime}$ and $n_{-}^{\prime}$ denote the number of new dialogues containing mental manipulation and not containing mental manipulation, respectively.
From Equation 2, EvoSA can expand nearly 5,000 high-quality dialogues from just 100 labeled ones. To further validate the generated data, we manually evaluated a random sample of the generated dialogues. Examples and evaluation results are provided in Appendix A.
3.2 Multi-Task Data Generation
Human teaching and learning are mutually reinforcing processes, where prior knowledge guides attention and deepens understanding. For large language models (LLMs), leveraging related-task knowledge improves learning efficiency and overall performance (Wang and Chang, 2022; Hsieh et al., 2023; Zhang et al., 2024; Chen et al., 2024; Yang et al., 2025a). Teaching and learning as complementary activities foster continuous growth and skill development (Fiorella and Mayer, 2013).
Inspired by this, our approach targets detecting mental manipulation in multi-turn dialogues, a complex task requiring subtle understanding and precise judgment. We design three interconnected tasks simulating a teaching–learning cycle to drive model improvement:
  • Task 1 (Feedback Generation, simulating teaching): The student generates feedback identifying errors in incorrect rationales.
  • Task 2 (Judgment with Rationale, simulating learning): The student provides binary judgments with rationale.
  • Task 3 (Judgment Only): The student provides binary judgments without explanation.
For Task 1, we first construct incorrect rationales as follows:
$r_{i}^{-}=\mathcal{T}(P_{ir}(x_{i}^{\prime},y_{i}^{\prime}))$ (3)
where $P_{ir}$ is a prompt template used to drive the teacher model to generate the incorrect rationale. $r_{i}^{-}$ is the incorrect rationale corresponding to the $i$-th sample $(x_{i}^{\prime},y_{i}^{\prime})$. Afterwards, we prompt the teacher model to generate feedback for incorrect answers, denoted as
$f_{i}=\mathcal{T}(P_{f}(x_{i}^{\prime},y_{i}^{\prime},r_{i}^{-}))$ (4)
where $P_{f}$ is a prompt template used to drive the teacher model to generate feedback on an incorrect rationale. $f_{i}$ is the feedback corresponding to the $i$-th incorrect rationale.
For Task 2, the correct rationale is generated by the teacher model according to the following expression:
$r_{i}^{+}=\mathcal{T}(P_{cr}(x_{i}^{\prime},y_{i}^{\prime}))$ (5)
where $r_{i}^{+}$ is the correct rationale corresponding to the $i$-th sample $(x_{i}^{\prime},y_{i}^{\prime})$, and $P_{cr}$ denotes the prompt template used to guide the teacher model in generating the correct rationale.
For Task 3, only the label is required. Prompt templates used for multi-task data generation are detailed in Appendix B.
3.3 Anti-Curriculum Distillation for Detecting Mental Manipulation
Anti-curriculum learning, which models human learning from hard to easy tasks, has proven effective despite seeming counterintuitive (Wu et al., 2021; Liu et al., 2022; Jarca et al., 2025). However, simultaneously learning multiple tasks with varying difficulty levels often leads to imbalanced training progress, which can negatively impact the performance of the primary task. To address this challenge and enhance the model’s effectiveness in detecting mental manipulation in multi-turn dialogues, a hard-to-easy distillation strategy, namely anti-curriculum distillation, is adopted as illustrated in Figure 2.
In the first stage, we prompt the student model $\mathcal{S}$ to act as a teacher by providing it with incorrect rationales without explicitly indicating the error, and asking it to identify the flaw and provide feedback. Mathematically, this can be formulated as
$f_{i}^{\mathcal{S}}=\mathcal{S}(P_{f}^{\mathcal{S}}(x_{i}^{\prime},y_{i}^{\prime},r_{i}^{-}))$ (6)
where $P_{f}^{\mathcal{S}}$ prompts the student to generate feedback, and $f_{i}^{\mathcal{S}}$ is the feedback on the incorrect rationale. We then define the following objective to train the student on Task 1:
$\mathcal{L}_{1}=\frac{1}{n\textquoteright}\sum_{i=1}^{n^{\prime}}[\ell(f_{i}^{\mathcal{S}},f_{i})]$ (7)
where $\ell(\cdot)$ is the cross-entropy loss function. The loss function for the first stage is
$\mathcal{L}^{s}_{1}=\mathcal{L}_{1}$ (8)
After the first stage, the student model gains an initial understanding of mental manipulation detection. In the second stage, it is further trained on both Task 1 and Task 2. For Task 2, the model must judge if a dialogue contains mental manipulation and provide a rationale, as formalized below:
$r_{i}^{\mathcal{S}}=\mathcal{S}(P_{r}^{\mathcal{S}}(x_{i}^{\prime},y_{i}^{\prime}))$ (9)
where $P_{r}^{\mathcal{S}}$ is a prompt template for guiding the student model to complete Task 2, and $r_{i}^{\mathcal{S}}$ is the judgement of the student model (with rationale). We first define the loss function for Task 2 as
$\mathcal{L}_{2}=\frac{1}{n\textquoteright}\sum_{i=1}^{n^{\prime}}[\ell(r_{i}^{\mathcal{S}},r_{i}^{+})]$ (10)
After that, we define the loss function for the second stage as
$\mathcal{L}^{s}_{2}=\mathcal{L}_{1}+\mathcal{L}_{2}$ (11)
For Task 3, the student model is required to provide a binary judgment, which can be formulated as follows:
$y_{i}^{\mathcal{S}}=\mathcal{S}(P_{y}^{\mathcal{S}}(x_{i}^{\prime}))$ (12)
where $P_{y}^{\mathcal{S}}$ is the prompt template guiding the student model’s binary judgment, and $y_{i}^{\mathcal{S}}$ denotes the student model’s binary output. The loss function for Task 3 is then defined as
$\mathcal{L}_{3}=\frac{1}{n\textquoteright}\sum_{i=1}^{n^{\prime}}[\ell(y_{i}^{\mathcal{S}},y_{i}^{\prime})]$ (13)
Subsequently, the loss function for the third stage is defined as
$\mathcal{L}^{s}_{3}=\mathcal{L}_{1}+\mathcal{L}_{2}+\mathcal{L}_{3}$ (14)
Overall, Task 1 is learned across all three stages, Task 2 across two stages, and Task 3 only in the final stage. This design ensures that more cognitively demanding tasks receive greater training exposure, thereby enhancing the model’s overall capability. By first engaging in Task 1 and Task 2, the student model learns to identify flaws in incorrect rationales and articulate correct reasoning. This foundation enables more effective learning of Task 3, which requires concise yet accurate judgments without explanatory support.
4 ReaMent Dataset
Although we propose EvoSA for unsupervised data synthesis, the field relies heavily on the single dataset MentalManip and lacks real-world data. To address this, we constructed ReaMent, derived from real-world multi-turn dialogues for detecting mental manipulation.
In our search, we found YTD-18M, a dataset of 18 million video-based dialogues from 20 million web videos, comprising four subsets: GPT-3-generated dialogues (Safe and Unsafe) for training converter models, and dialogues generated by these models to train CHAMPAGNE (Safe and Unsafe). We selected GPT-3-generated safe and unsafe dialogues and unsafe converter-generated dialogues, totaling 102,323 samples.
To reduce human annotation costs, we used the Qwen2.5-3B-Instruct (Qwen Team, 2024) distilled via MentalMAC, identifying 9,401 dialogues potentially containing mental manipulation from 102,323 samples. After filtering out low-quality data, twelve annotators labeled each dialogue, with three annotators assigned per sample. We curated ReaMent, a dataset of 5,000 high-quality samples, separated into subsets based on consensus agreement (ReaMentcon) and majority agreement (ReaMentmaj). Annotation details 390 and statistics are in Appendix C.
5 Experiments
Model Method MentalManip ReaMent
Acc Pre Re F1mac F1w Acc Pre Re F1mac F1w
DeepSeek-R1 0.739 0.730 0.988 0.571 0.674 0.665 0.663 0.952 0.534 0.598
GPT-4o 0.739* 0.741* 0.952* 0.617* 0.620 0.674 0.766 0.570 0.608
GPT-4 0.677* 0.813* 0.691* 0.649* 0.687* 0.593 0.676 0.676 0.564 0.593
GPT-4 Turbo 0.657* 0.788* 0.682* 0.629* 0.651 0.683 0.827 0.589 0.630
Claude-3.5-Sonnet 0.734 0.778 0.861 0.664 0.723 0.582 0.702 0.581 0.573 0.589
Llama3-70B-Instruct 0.720 0.761 0.868 0.637 0.703 0.611 0.631 0.916 0.451 0.527
Qwen2.5-3B Vanilla 0.624 0.793 0.618 0.603 0.638 0.578 0.765 0.474 0.578 0.580
CoT 0.396 0.859 0.151 0.374 0.330 0.390 0.769 0.042 0.312 0.252
IAP 0.461 0.835 0.275 0.458 0.441 0.409 0.679 0.116 0.365 0.322
SFT 0.686 0.743 0.834 0.599 0.670 0.643 0.700 0.756 0.606 0.637
Distill 0.692 0.796 0.746 0.653 0.697 0.629 0.716 0.664 0.615 0.633
Ours 0.758 0.771 0.925 0.667 0.734 0.713 0.716 0.899 0.653 0.690
Table 1: Performance of large LLMs and student models on MentalManip and ReaMent. Acc = Accuracy; Pre = Precision; Re = Recall; F1mac/w = macro- and weighted-F1. * indicates data from original papers. "Claude-3.5-Sonnet" and "Qwen2.5-3B" refer to Claude-3.5-Sonnet-20241022 and Qwen2.5-3B-Instruct, respectively. "IAP" and "Distill" stand for intent-aware prompting and distilling step-by-step. Llama3-70B-Instruct is the sole teacher model used. The best and second-best student model results are bolded and underlined, respectively.
[…]
3B Vanilla 0.624 0.793 0.618 0.603 0.638
Ours 0.758 0.771 0.925 0.667 0.734
$\Delta$ (%) +21.5 -2.8 +49.7 +10.6 +15.1
1.5B Vanilla 0.422 0.837 0.203 0.410 0.379
Ours 0.707 0.772 0.816 0.643 0.701
$\Delta$ (%) +67.5 -7.8 +302.0 +56.8 +85.0
0.5B Vanilla 0.484 0.670 0.499 0.461 0.503
Ours 0.664 0.735 0.805 0.581 0.652
$\Delta$ (%) +37.2 +9.7 +61.3 +26.0 +29.6
Table 2: Performance of student models with different sizes on MentalManip. Acc = Accuracy; Pre = Precision; Re = Recall; F1mac/w = macro- and weighted-F1.
[…]
3B Vanilla 0.624 0.793 0.618 0.603 0.638
Ours 0.758 0.771 0.925 0.667 0.734
$\Delta$ (%) +23.4 -6.4 +89.7 +13.0 +19.0
1.5B Vanilla 0.422 0.837 0.203 0.410 0.379
Ours 0.707 0.772 0.816 0.643 0.701
$\Delta$ (%) +44.6 -4.0 +344.0 +47.3 +69.3
0.5B Vanilla 0.484 0.670 0.499 0.461 0.503
Ours 0.664 0.735 0.805 0.581 0.652
$\Delta$ (%) +17.0 -2.3 +88.3 +7.9 +16.5
Table 3: Performance of student models with different sizes on ReaMent. Acc = Accuracy; Pre = Precision; Re = Recall; F1mac/w = macro- and weighted-F1.
5.1 Experiment Setting
Dataset and Models. Given the subtle and deceptive nature of mental manipulation, we employ MentalManipcon [^5] Labels based on unanimous agreement among three annotators and ReaMentcon, with 1,749 and 3,782 dialogues respectively, ensuring accurate evaluation labels. Both datasets are stratified and split 6:2:2 into training, validation, and test sets. We use LLaMA3-70B-Instruct (AI@Meta, 2024) as the teacher model. For the main experiment, we adopt Qwen2.5-3B-Instruct as the student model, and additionally use Qwen2.5-1.5B-Instruct and Qwen2.5-0.5B-Instruct to examine the impact of model size on MentalMAC (Qwen Team, 2024).
Baselines. We evaluate five categories: (1) SOTA LLMs, including DeepSeek-R1 (Guo et al., 2025), GPT-4o (Hurst et al., 2024), GPT-4 (Achiam et al., 2023), GPT-4 Turbo (OpenAI, 2023), Claude-3.5-Sonnet-20241022 (Anthropic, 2024), and Llama3-70B-Instruct (AI@Meta, 2024); (2) vanilla student models; (3) existing techniques for enhancing LLM performance on this task, including CoT (Yang et al., 2024) and intent-aware (Ma et al., 2025) prompting; (4) a general-purpose distillation method, namely distilling step-by-step (Hsieh et al., 2023), which has been validated to outperform standard KD; and (5) supervised fine-tuning (Yang et al., 2024).
Implementation Details. All experiments were conducted on two NVIDIA H800 GPUs using the Lion optimizer (Chen et al., 2023) and LoRA (Hu et al., 2022). Training was performed for 6 epochs on MentalManip and 5 epochs on ReaMent, with LoRA configured at rank 8, a learning rate of 5e-6, batch size of 8, and gradient accumulation set to 4. During inference, we used a temperature of 0.1 and top-p of 0.5. Additional implementation details can be found in Appendix D.
5.2 Results Analysis
Method MentalManip ReaMent
Acc Pre Re F1mac F1w Acc Pre Re F1mac F1w
w/o EvoSA 0.714 0.774 0.830 0.645 0.706 0.670 0.717 0.785 0.634 0.663
w/o Task 1 0.682 0.783 0.743 0.639 0.686 0.667 0.707 0.802 0.623 0.656
w/o Task 2 0.719 0.790 0.808 0.666 0.716 0.652 0.721 0.728 0.626 0.651
w/o Task 3 0.744 0.767 0.906 0.656 0.723 0.692 0.715 0.848 0.642 0.676
w/o AC (joint) 0.713 0.749 0.877 0.621 0.691 0.661 0.740 0.721 0.637 0.662
w/o AC (cur) 0.650 0.820 0.633 0.631 0.663 0.626 0.732 0.639 0.615 0.632
Ours 0.758 0.771 0.925 0.667 0.734 0.713 0.716 0.899 0.653 0.690
Table 4: Results of the ablation study. Please note that we used Qwen2.5-3B-Instruct for the ablation study. Acc = Accuracy, Pre = Precision, Re = Recall, F1mac/w = macro- and weighted-F1 scores. "w/o" indicates the exclusion of a specific module. "w/o AC (joint)" means not using anti-curriculum learning, but traditional joint learning. "w/o AC (cur)" indicates that instead of anti-curriculum learning, curriculum learning is applied, meaning the tasks are learned in the order from Task 3 to Task 1. The best performance is marked in bold.
MentalMAC consistently outperforms all baselines on both MentalManip and ReaMent datasets, as shown in Table 1. Compared to the strongest student baseline, it achieves notable gains: accuracy improves by 9.54% and 10.89%, recall by 10.91% and 18.92%, macro F1 by 2.14% and 5.82%, and weighted F1 by 5.31% and 8.32% on MentalManip and ReaMent, respectively. Crucially, MentalMAC enables the 3B-parameter Qwen2.5-3B-Instruct to surpass large-scale LLMs like DeepSeek-R1 and GPT-4o on key metrics, proving our method’s effectiveness and efficiency.
To evaluate performance across model sizes, we tested Qwen2.5-1.5B-Instruct and Qwen2.5-0.5B-Instruct as well. Tables 2 and 3 present results on MentalManip and ReaMent, respectively. On MentalManip, recall increases by 49.7% for the 3B model and by 61.3% even for the smallest 0.5B model; on ReaMent, the 1.5B model’s recall improves by 344.0%.
Although precision slightly declines, substantial recall gains consistently boost accuracy, macro F1, and weighted F1, enhancing overall performance and sensitivity to subtle mental manipulation across model scales.
5.3 Ablation Study
We analyze the ablation results in Table 4 to assess the effectiveness of key components. Removing EvoSA slightly improves precision but significantly reduces recall, which leads to overall performance declines and highlights EvoSA’s role in capturing diverse manipulative patterns. Excluding any of the three tasks causes performance to drop, with Tasks 1 and 2 being more important than Task 3, likely because of implicit learning through task interrelations. Anti-curriculum learning is also more effective than joint or curriculum learning, as tackling the more difficult Tasks 1 and 2 first results in better generalization on Task 3.
5.4 Cross-Dataset Generalization Analysis
Setting Acc Pre Re F1mac F1w
[…]
To assess MentalMAC’s generalization across different data distributions, we performed a cross-dataset evaluation using Qwen2.5-3B-Instruct. As shown in Table 5, although the model’s performance slightly decreases when tested on a different dataset, it remains consistently strong, demonstrating robust generalization capability.
Interestingly, we observe that the model trained on ReaMent achieves better results on MentalManip than on ReaMent itself. This may be attributed to the nature of MentalManip’s movie-based dialogues, which tend to be more dramatized yet simpler, while ReaMent contains richer and more complex real-world conversations. Consequently, training on ReaMent enables the model to capture more generalizable features.
6 Conclusion
We propose MentalMAC, a novel method that enhances LLMs’s ability to detect mental manipulation in multi-turn dialogues via multi-task anti-curriculum distillation. To overcome the challenge of limited training data, we introduce EvoSA, an unsupervised data expansion strategy grounded in evolutionary operations and speech act theory. Our approach defines three progressively simpler tasks: providing feedback on incorrect rationales, generating correct answers with rationales, and producing answers only. A teacher model generates training samples for each task, while the student model is trained in a hard-to-easy order. Leveraging the distilled student model, we support human annotation efforts and construct ReaMent, a dataset of 5,000 real-world dialogues that addresses a critical gap in mental manipulation research. Future work will explore more efficient distillation strategies to further enhance model performance.
Limitations
We recognize the following limitations of the work in this paper:
MentalMAC Although MentalMAC significantly narrows the gap between student and teacher models and outperforms the strongest baselines, it incurs a slight precision decrease that may entail therapeutic costs. Future work could explore improving detection accuracy without sacrificing precision. Additionally, the order of task learning was predetermined in this study, but when to learn each task to maximize model performance remains to be further investigated.
EvoSA Our analysis of the data synthesized by EvoSA (see Appendix A for details) reveals that the quality of generated samples may degrade when the parent dialogue is too short, indicating a dependency on sufficient contextual input. Furthermore, we did not explore how EvoSA performs across different LLMs, leaving open the question of model-specific effectiveness. Lastly, due to the safety constraints imposed by LLMs, EvoSA is currently unable to generate dialogues that simultaneously contain toxic content and elements of mental manipulation.
ReaMent Real-world dialogues are often multi-party, whereas our ReaMent dataset contains only two-person conversations. Although we found that EvoSA can extend two-person dialogues into multi-party ones (see Appendix A for details), this remains a limitation of the ReaMent dataset. Additionally, ReaMent is currently limited to English content, which may restrict its applicability in multilingual or cross-cultural contexts.
Ethics Statement
Our data collection process was approved by the IRB. Before annotation, we identified that some dialogues contained toxic content, including sexually explicit and violent material. We informed annotators and evaluators beforehand about the possibility of encountering such content and obtained their informed consent. Additionally, we offer psychological support to annotators and evaluators experiencing distress during the annotation process. The data collected is publicly available and does not infringe upon the rights of others, and its release respects individual privacy. In addition, we took steps to avoid presenting highly disturbing dialogues in the paper, insofar as doing so would not compromise the integrity of our analysis.
Appendix A Evaluation of the Proposed EvoSA
Figure 4: An example of dialogue synthesized using EvoSA. Colour blocks of the same colour indicate parts of the parent dialogue that the child dialogue may potentially draw upon.
A.1 Example Demonstration
Our proposed EvoSA is a prompting method for synthesizing new dialogues by combining two parent dialogues into a single child dialogue. An example of such synthesis from MentalManip is shown in Figure 4. The blocks of the same colour represent portions of the child dialogue that closely resemble the parent dialogue. As observed, similar to the evolutionary algorithm, the child dialogue incorporates elements from the parent dialogue. Furthermore, the child dialogues exhibit a rich variety of speech acts, distinguishing them from the parent dialogues in terms of background, roles, and content. In addition, as illustrated in Figure 5, we found that EvoSA can extend a two-person dialogue to a multi-person conversation, demonstrating its potential to synthesize dialogues involving multiple participants.
In data synthesis, we did not find refusals to answer on MentalManip, probably because MentalManip does not contain excessively toxic text. However, during the expansion of ReaMent, we observed that the LLM declined to respond in 257 out of 3,257 instances as a result of its safety mechanisms. Although there is less than a 10% chance of refusal to respond, this does not significantly affect our ability to generate sufficient data, provided we have a certain number of parent dialogues.
Figure 5: An example of a multi-person dialogue generated by EvoSA. Colour blocks of the same colour indicate parts of the parent dialogue that the child dialogue may potentially draw upon.
A.2 Quality of Child Dialogues
In selecting evaluators, we recruited candidates via a voluntary application process. They were informed about the potential exposure to offensive content and provided their consent. Subsequently, the candidates received training on mental manipulation. After that, we randomly selected 100 dialogues from MentalManipcon for candidates to annotate. Only those with 80% annotation accuracy were considered as evaluators. In the end, we recruited 3 evaluators to evaluate the quality and label consistency of the child dialogues.
The evaluators were asked to rate the quality of each dialogue on a five-point scale (from 1 to 5, low to high), considering aspects such as fluency and logical coherence. The instruction we provided to the evaluators is:
Please rate the quality of this dialogue on a 5-point scale (1 = very poor, 5 = excellent), taking into account factors such as coherence, logical consistency, and naturalness.
child Dialogue:
<insert child dialogue>
We collected the scores assigned by the three evaluators to the Yes and No labels, and reported their average values in Table 6. Analysis shows that child dialogues labeled No tend to be of lower quality than those labeled Yes. On average, child dialogues labeled Yes and No received scores of 4.265 and 3.788, respectively. The overall mean score of 4.022 indicates that EvoSA effectively guides LLMs to generate higher-quality dialogues. Figure 6 illustrates the correlation between dialogue length and quality score. The regression line shows that as dialogue length increases, quality tends to improve, likely due to the additional context provided by longer dialogues. The Pearson correlation coefficient (r = 0.270, p = 0.0004) indicates a statistically significant positive correlation, suggesting that dialogue length does have an effect on the quality score in EvoSA-generated dialogues.
     Setting      Yes      No      Mean
     E1      4.357      3.794      4.070
     E2      4.337      3.922      4.125
     E3      4.102      3.647      3.870
     Mean      4.265      3.788      4.022
Table 6: Results of the quality assessment for child dialogues generated by EvoSA. "E1", "E2", and "E3" refer to the three evaluators we recruited, while "Yes" and "No" represent the dialogue labels.
Figure 6: Relationship between dialogue quality and dialogue length. The Pearson correlation coefficient (r = 0.270, p = 0.0004) indicates a statistically significant positive correlation. Blue circles represent individual dialogues, with length on the x-axis and average quality score on the y-axis. The red dashed line shows the regression trend, with the shaded area representing the 95% confidence interval.
A.3 Label Consistency Between Child and Parent Dialogues
To assess label consistency, evaluators were asked to analyze why parent dialogues were labeled as containing or not containing mental manipulation. Based on their analysis, they then determined whether the labels of the child dialogues were consistent with those of the parent dialogues. We provided the following instruction for evaluators to complete the task:
Please start by analyzing why both Dialogue 1 and Dialogue 2 (the parent dialogues) are labeled <Yes/No>. Then, based on your analysis, determine whether the label of the child dialogue is consistent with the parent dialogues. Mark 1 if consistent; otherwise, mark 0.
Dialogue 1:
<insert dialogue 1>
Dialogue 2:
<insert dialogue 2>
Child Dialogue:
<insert child dialogue>
Table 7 presents the results of the labeling consistency assessment. When two out of three evaluators consider the labels consistent ("$\geq 2$"), we regard the child dialogue as consistent with the parent dialogues in terms of label. Of the child dialogues labeled No and Yes, 89.2% and 96.9%, respectively, are being consistent with the parent dialogue labels. This indicates that dialogues without mental manipulation are being rated lower than those containing mental manipulation, likely because LLMs are tending to mistakenly classify non-manipulative dialogues as manipulative.
Setting Yes No Total
$<2$ 3 (3.1%) 11 (10.8%) 14 (7.0%)
$\geq 2$ 95 (96.9%) 91 (89.2%) 186 (93.0%)
Total 98 102 200
Table 7: Results of the label consistency assessment. "$<2$" and "$\geq 2$" denote that fewer than 2 and at least 2 people, respectively, believe the labels are consistent.
To assess the relationship between label consistency, dialogue length, and dialogue quality, we define the following consistency scores:
$c_{i}=\frac{1}{n_{e}}\sum_{j=1}^{n_{e}}e_{ji}$ (15)
where $c_{i}$ denotes the label consistency score of the $i$-th child dialogue; $e_{ji}$ denotes the $j$-th evaluator’s rating of the $i$-th child dialogue in terms of label consistency, taking the value 1 if consistent and 0 otherwise; and $n_{e}$ denotes the number of evaluators. Figures 7 and 8 illustrate the relationships between label consistency scores and dialogue length, and dialogue quality, respectively. The Pearson correlation coefficient indicates that there is no significant relationship between label consistency scores and dialogue length, but it shows a significant positive correlation with dialogue quality.
Figure 7: Relationship between label consistency score and dialogue length. The Pearson correlation coefficient (r = -0.060, p = 0.4025) indicates a weak negative correlation, which is not statistically significant. Blue circles represent individual dialogues, with length on the x-axis and label consistency score on the y-axis. The red dashed line shows the regression trend, with the shaded area representing the 95% confidence interval.
Figure 8: Relationship between label consistency score and dialogue quality. The Pearson correlation coefficient (r = 0.247, p = 0.0004) indicates a statistically significant positive correlation. Blue circles represent individual dialogues, with average quality score on the x-axis and label consistency score on the y-axis. The red dashed line shows the regression trend, with the shaded area representing the 95% confidence interval.
Appendix B Prompting Formats
B.1 Prompt Templates for Teacher Models
The prompt used to guide the teacher model in generating both correct and incorrect responses with rationales is the same:
You are an advanced dialogue analysis agent.
Using your knowledge of dark psychology based on the definition of mental manipulation, please explain why this dialogue <does not contain/contains> elements of mental manipulation. Let’s think step by step.
### Definition of Mental Manipulation:
Mental manipulation is using language to influence, alter, or control an individual’s psychological state or perception for the manipulator’s benefit.
### Dialogue:
<insert dialogue>
### Output Format:
Rationale: [Provide only strong evidence using direct dialogue quotes. Clearly explain how the language used aligns—or does not align—with known manipulation tactics.]
where <does not contain/contains> is given based on true labels. The definition of mental manipulation is adopted from Wang et al. (2024). As can be seen, we require the teacher model not only to reason step by step, but also to provide detailed evidence explaining why the dialogue does or does not conform to known manipulation tactics. In order to generate incorrect answers, here we prompt the teacher model with incorrect labels.
We used the following prompt template to allow the teacher model to give feedback on incorrect responses that contained rationales:
You are an advanced dialogue analysis teacher.
In the task of detecting whether the dialogue contained elements of mental manipulation, students gave incorrect answers. You should point out the mistakes in the student’s answer using knowledge of dark psychology based on the definition of mental manipulation. Let’s think step by step.
### Definition of Mental Manipulation:
Mental manipulation is using language to influence, alter, or control an individual’s psychological state or perception for the manipulator’s benefit.
### Dialogue:
<insert dialogue>
### Student’s Answer:
This dialogue <does not contain/contains> elements of mental manipulation. <insert incorrect response>
### Hint:
This dialogue <does not contain/contains> elements of mental manipulation.
### Output Format:
Feedback: [Provide student’s mistakes.]
where <does not contain/contains> is given based on true labels.
B.2 Prompt Templates for Student Models
Obviously, the student model prompt could not mirror the teacher model prompt exactly; therefore, we designed a simplified version specifically for the student. The prompt template for Task 1 is
In the task of detecting whether the dialogue contains elements of mental manipulation, one student gave an answer. You should point out any mistakes (if any) in the student’s response.
### Dialogue:
<insert dialogue>
### Student’s Response:
<insert incorrect response>
The prompt template for Task 2 is
I will provide you with a dialogue. Your task is to analyse whether the dialogue contains elements of mental manipulation. Tell me your judgement and rationale.
<insert dialogue>
The prompt template for Task 2 makes the student model first output a binary judgment (Yes or No). To save computational resources, we use the same prompt for both Task 1 and Task 2. For Task 1, we compute the loss solely on the token corresponding to the binary judgement.
Appendix C Details of Our ReaMent Dataset
C.1 Details on the Annotation Process
In selecting annotators, we recruited candidates via a voluntary application process. They were informed about the potential exposure to offensive content and provided their consent. Subsequently, the candidates received training on mental manipulation. After that, we randomly selected 100 dialogues from MentalManipcon for candidates to annotate. Only those with 80% annotation accuracy were considered as annotators. In the end, we recruited 12 annotators and divided them into groups of three each annotating the same data.
To ensure the quality of the annotation, we had the annotator learn about mental manipulation beforehand. After that, we provided them with the initially screened data along with the following annotation guidelines:
Based on the definition of mental manipulation, please determine whether the given dialogue contains elements of mental manipulation. If it does, label it as 1; otherwise, label it as 0. In addition, provide your confidence level in the annotation on a 5-point scale (1 = very uncertain, 5 = very confident).
### Definition:
Mental manipulation is using language to influence, alter, or control an individual’s psychological state or perception for the manipulator’s benefit.
### Dialogue:
<insert dialogue>
where the definition of mental manipulation comes from Wang et al. (2024). Following this instruction, we collected both the labels assigned by all annotators and their corresponding confidence levels.
C.2 Detailed Descriptions
ReaMent contains annotations indicating whether each dialogue includes elements of mental manipulation, labeled as either Yes or No. For ReaMent, a dialogue is labeled as Yes if at least two out of the three annotators agree on that label; otherwise, it is labeled as No. We denote the dialogues from ReaMent where all three annotators agreed on the label as ReaMentcon, and those where there was majority agreement among the annotators as ReaMentmaj.
Table 8 presents the distribution of binary labels. As shown, the number of dialogues labeled as No is slightly smaller than those labeled as Yes.
Dataset Yes No Y/N
ReaMentcon 2,368 1,414 1.67:1
ReaMentmaj 2,813 2,187 1.29:1
Table 8: Distribution of binary labels on the ReaMent dataset. "Yes" and "No" indicate the number of dialogues labeled as such, while "Y/N" denotes their ratio.
C.3 Analysis of Annotation Quality
For each annotator group, we evaluated annotation quality using inter-annotator agreement scores as defined by Wang et al. (2024):
$s_{ij}=\frac{1}{n_{a}}\sum_{k=1}^{n_{c}}I(a_{ik},a_{jk})$ (16)
where $s_{ij}$ represents the agreement score between the $i$-th and $j$-th annotators, and $n_{c}$ donates the size of the co-annotated data. $a_{ik}$ is the annotation result of the $i$-th annotator on the $k$-th data point. The indicator function $I(\cdot,\cdot)$ takes the value 1 when $a_{ik}$ and $a_{jk}$ are the same, and 0 otherwise. Figure 9 presents a heatmap of the consistency scores, with values ranging from 0.67 to 0.95, indicating good data quality.
Figure 9: Inter-annotator agreement. GiAj represents the j-th annotator of group i.
To further evaluate the reliability of labels, we define a label confidence score as follows:
$v_{k}=\frac{\sum_{i=1}^{n_{a}}c_{ik}I(a_{ik},t_{k})}{\sum_{i=1}^{n_{a}}c_{ik}}$ (17)
where $v_{k}$ denotes the label confidence score for the $k$-th data instance. $n_{a}$ is the number of annotations for the same instance, and $c_{ik}$ is the confidence level of the $i$-th annotator in their annotation of the $k$-th instance (in this study, specified on a scale from 1 to 5). $t_{k}$ represents the final assigned label.
It can be analyzed that the label confidence score ranges from 0 to 1, with higher values indicating greater reliability of the final label. Figure 10 shows the distribution of label confidence scores. The mean, first quartile, and median suggest relatively high label credibility.
Figure 10: Distribution of label confidence scores. "Mean" refers to the average value, while "Q1" and "Q2" is the first quartile and the median, respectively.
Appendix D Experimental Parameters
D.1 Parameters for Teacher Model
The teacher model performs two tasks: the first involves the unsupervised synthesis of new dialogues using EvoSA, and the second concerns the construction of multitask data. For both tasks, we use identical parameters, setting the temperature to 0 and the maximum token limit to 512.
Furthermore, we employed EvoSA to perform unsupervised data augmentation on the training sets of MentalManipcon and ReaMentcon, generating an additional 3,000 and 2,000 samples, respectively, with a balanced 1:1 ratio of the binary labels.
D.2 Parameters for Student Model
During training, the student model is trained with a maximum sequence length of 1500 tokens. In the first two stages we train 2 epochs, and in the last stage we train 2 epochs on MentalManip and 1 epoch on ReaMent. In each stage, learning rate warm-up with a constant schedule is applied, with the warm-up ratio set to 0.05.
During the inference phase, to ensure the output contains valid information, we set the maximum number of tokens to 5, enable sampling by setting do_sample to True, and configure the temperature and top-p parameters to 0.1 and 0.5, respectively.